# Cert-SSB: Toward Certified Sample-Specific Backdoor Defense

### v1

@@ Abstract
Deep neural networks (DNNs) are vulnerable to backdoor attacks, where an attacker manipulates a small portion of the training data to implant hidden backdoors into the model. The compromised model behaves normally on clean samples but misclassifies backdoored samples into the attacker-specified target class, posing a significant threat to real-world DNN applications. Currently, several empirical defense methods have been proposed to mitigate backdoor attacks, but they are often bypassed by more advanced backdoor techniques. In contrast, certified defenses based on randomized smoothing have shown promise by adding random noise to training and testing samples to counteract backdoor attacks. In this paper, we reveal that existing randomized smoothing defenses implicitly assume that all samples are equidistant from the decision boundary. However, it may not hold in practice, leading to suboptimal certification performance. To address this issue, we propose a sample-specific certified backdoor defense method, termed Cert-SSB. Cert-SSB first employs stochastic gradient ascent to optimize the noise magnitude for each sample, ensuring a sample-specific noise level that is then applied to multiple poisoned training sets to retrain several smoothed models. After that, Cert-SSB aggregates the predictions of multiple smoothed models to generate the final robust prediction. In particular, in this case, existing certification methods become inapplicable since the optimized noise varies across different samples. To conquer this challenge, we introduce a storage-update-based certification method, which dynamically adjusts each sample’s certification region to improve certification performance. We conduct extensive experiments on multiple benchmark datasets, demonstrating the effectiveness of our proposed method. Our code is available at https://github.com/NcepuQiaoTing/Cert-SSB.
Index Terms:
Certified Backdoor Defense, Backdoor Defense, Randomized Smoothing, Trustworthy ML, AI Security
@@ I Introduction
Figure 1: An overview of existing randomized smoothing-based certified backdoor defenses and our Cert-SSB. The existing methods apply fixed noise to smooth classifiers for all inputs, ignoring sample diversity. This often leads to suboptimal certification performance. In contrast, Cert-SSB optimizes the noise, enabling the smoothing strategy to adapt to different inputs (as shown in the right figure), thereby achieving more robust certified backdoor defenses.
@@
To mitigate backdoor threats, researchers have made significant efforts in both detection [10, 11, 12] and defense [13, 14, 15]. However, advanced backdoor attacks [16, 17, 18] can still easily bypass existing defenses, leading to an ongoing arms race between defenders and attackers. To address this issue, some studies have proposed certified backdoor defense methods, primarily categorized into deterministic certification [19, 20, 21, 22] and probabilistic certification [23, 24]. These methods aim to provide theoretical guarantees, ensuring that the classification results of testing samples remain consistent regardless of whether the model is trained on clean or backdoor data, as long as the perturbation induced by the trigger remains within an $\ell_{p}$ norm ball of radius $r$. However, deterministic methods face scalability challenges when applied to large-scale neural networks. Consequently, probabilistic certification approaches based on randomized smoothing have emerged as a more practical alternative and have demonstrated robustness on large-scale datasets such as ImageNet [25]. Randomized smoothing was initially developed to certify robustness against adversarial examples. Its principle is to introduce random noise into the input data, ensuring that the classification results remain consistent within a specified region ($e.g.$, an $\ell_{p}$ norm neighborhood), thereby achieving robustness. Notably, pioneering studies [23, 24] showed that certified backdoor defenses based on random smoothing, which are robust against bounded backdoor patterns ($i.e.$, constrained pixel-level perturbation), can be achieved by introducing isotropic Gaussian noise into a tuple consisting of a testing instance and the training set to mitigate the impact of attacker-injected triggers, effectively neutralizing backdoor attacks during the training phase.
@@
Fortunately, the answer to the above question is affirmative. Arguably, the most direct approach is to optimize the noise at each sample by maximizing the confidence margin between the top-1 and top-2 predicted classes of the classifier ($i.e.$, the certification radius). However, due to the lack of an analytical expression for the certification radius, direct optimization is challenging. Inspired by [27], we employ a stochastic gradient ascent method to iteratively optimize the noise in order to maximize the certification radius. However, during the optimization process, the continuous adjustment of noise alters the data distribution, leading to increased variance in gradient estimation and affecting optimization stability. To address this issue, we propose an advanced sample-specific certified backdoor defense method, termed Cert-SSB. In general, Cert-SSB consists of two main stages: training and inference. In the first stage, we train multiple smoothed models using the optimized noise, which is obtained through stochastic gradient ascent to maximize the certification radius. Generally, the certification radius is computed based on the predictions of classifiers trained with fixed noise. Besides, we adopt a reparameterization technique to reduce gradient variance and enhance optimization stability. In the inference stage, we aggregate multiple smoothed classifiers trained in the first stage to generate the final smoothed prediction. However, since the optimized noise results in different noise magnitudes for each sample, existing certification methods, which typically assume a fixed noise level, are no longer directly applicable. To resolve this issue, we propose a storage-update-based certification method, which dynamically adjusts the certification region ($i.e.$, the space covered by the certification radius) for each sample. This ensures that certification regions do not overlap between different samples and that predictions remain consistent within each certified region.
@@
  • We propose a sample-specific certified backdoor defense method ($i.e.$, Cert-SSB) to dynamically adjust the smoothing noise magnitude for different samples to optimize certification performance.
@@
  • We conduct extensive experiments on benchmark datasets to validate Cert-SSB’s effectiveness, demonstrating its superior certification performance over existing methods.
The remainder of this paper is organized as follows. Section II provides a brief overview of related work. Section III revisits the limitations of existing certified backdoor defenses. Following that, in Section IV, we describe the threat model, defense goal, and an overview of our approach, followed by a detailed design of our certified backdoor defense method. Section V reports experimental results and performance analysis. Finally, we conclude this paper at the end.
@@ II-A Backdoor Attacks
Backdoor attacks [28, 29, 30, 31] have emerged as a new threat in the training process of deep neural networks (DNNs). There are various ways to categorize backdoor attacks. First, they can be classified based on the adversarial objective into all-to-one and all-to-all. In all-to-one attacks, all samples with backdoor triggers are misclassified into a predefined fixed target label, making the attack relatively simple and straightforward. In contrast, all-to-all attacks involve misclassifying samples with triggers into specific target classes based on their original categories, with a fixed mapping relationship between classes. This type of attack is more complex. Second, backdoor attacks can be categorized based on the threat scenario into three major types: (1) poison-only attacks [32, 16], (2) training-controlled attacks [33, 34], and (3) model-modified attacks [35, 36]. Specifically, poison-only attacks restrict the adversary to modifying the training dataset; training-controlled attacks allow the adversary to fully control the training process, including both the training data and algorithms. In contrast to these approaches, model-modified attacks mainly focus on the deployment phase rather than the training phase, embedding hidden backdoors by directly modifying model weights or introducing malicious DNN modules. In this paper, we mainly focus on poison-only backdoor attacks, which represent the most classical setting and pose the broadest threat scenarios. Recently, there are also a few works exploring how to exploit backdoor attacks for positive purposes [37, 38, 39, 40, 41, 42], which is out of the scope of this paper.
@@ II-B2 Certified defenses
Existing certified defense methods can be categorized into deterministic defenses [19, 20, 21, 22], which provide a guaranteed ‘certified’ outcome when the input is robust to attacks, and probabilistic defenses [23, 24], which ensure a ‘certified’ result with a certain probability ($e.g.$, 99.9%), where the randomness is independent of the input sample. However, deterministic defenses often face scalability challenges when applied to large networks. In this work, we focus on probabilistic certified defenses.
Probabilistic certification offers better scalability. Previous methods primarily relied on intrinsic mechanisms [57, 58] or randomized smoothing techniques [23, 24] to achieve robust predictions. For example, Jia et al. [57] leveraged the intrinsic ensemble technique within the bagging mechanism, focusing on the number of poisoned samples. Similarly, Jia et al. [58] employed the intrinsic majority voting mechanism in the $k$-nearest neighbors algorithm, focusing on the number of neighbors. However, these methods do not consider the size of the trigger, making them unsuitable for defending against backdoor attacks. On the other hand, Wang et al. [23] were the first to apply randomized smoothing to defend against backdoor attacks, introducing random noise to neutralize potential triggers. However, this method lacked comprehensive experimental evaluation and failed to achieve a high robustness bound. Recently, Weber et al. [24] proposed the robustness against backdoor attacks (RAB) framework, which provides certified robustness against backdoor attacks and sets a theoretical benchmark for provable defenses in this field. The detailed implementation of RAB will be discussed in Section II-C. However, current methods implicitly assume that all samples are equidistant from the decision boundary, which may not hold in practice. This leads to suboptimal certification and highlights the urgent need for adaptive approaches that account for sample-specific characteristics.
@@ II-C Randomized smoothing and RAB
$g(\bm{x},\sigma):=\arg\max_{y\in\mathcal{Y}}\mathcal{P}_{\bm{\epsilon}}(f(\bm{x}+\bm{\epsilon})=y),$ (1)
where $\bm{\epsilon}\sim\mathcal{N}(0,\sigma^{2}I)$. The noise level $\sigma$ is a hyperparameter that controls the trade-off between robustness; it does not change with the input $\bm{x}$. Using the Neyman-Pearson lemma [60], Cohen et al. [59] proved that $g(\bm{x},\sigma)$ is certifiably robust to adversarial perturbations under the $\ell_{2}$ norm constraint. Define $y_{A}=\arg\max_{y}\mathcal{P}_{\bm{\epsilon}}(f(\bm{x}+\bm{\epsilon})=y)$, and assume that when classifying a perturbed input $\bm{x}+\bm{\epsilon}$, the base classifier $f$ assigns the most probable class $y_{A}$ with probability $P_{A}=\mathcal{P}_{\bm{\epsilon}}(f(\bm{x}+\bm{\epsilon})=y_{A})$, and the second most probable class $y$ with probability $P_{B}=\max_{y\neq y_{A}}\mathcal{P}_{\bm{\epsilon}}(f(\bm{x}+\bm{\epsilon})=y)$. Then, it is always true that $g(\bm{x}+\bm{\Delta},\sigma)=y_{A}$ as long as $\|\bm{\Delta}\|_{2}<r$, where the certified robust radius $r$ is given by:
$r(\bm{x},\sigma):=\frac{\sigma}{2}\left(\Phi^{-1}(P_{A}(\bm{x},\sigma))-\Phi^{-1}(P_{B}(\bm{x},\sigma))\right),$ (2)
@@
Figure 2: Distribution of $\ell_{2}$ norm distances between samples and their closest boundary samples.
@@
Overview of RAB [24]. Given a dataset $\mathcal{D}$ and a testing instance $\bm{x}$, the base classifier $f$, learns a probability distribution $\mathcal{P}_{\bm{\epsilon}}(\bm{q}(\bm{x},\mathcal{D})=y)$ over the class labels. It outputs the label that is most likely under this learned distribution $\bm{q}$: $f(\bm{x},\mathcal{D})=\arg\max_{y}\mathcal{P}_{\bm{\epsilon}}(\bm{q}(\bm{x},\mathcal{D})=y)$. A smoothed classifier $g(\bm{x},\mathcal{D},\sigma)$ returns whichever class the base classifer $f(\bm{x},\mathcal{D})$ is most likely to predict when $\bm{x}$ is perturbed by smoothing distributions $X=(Z,D)$:
@@
where $Z\sim\mathcal{N}(0,\sigma^{2}I)$ is assumed to be independent, and $D\sim\mathcal{N}(0,\sigma^{2}I)$ consists of $n$ independent and identically distributed random variables $D^{(i)}$, each added to a training instance in $\mathcal{D}$. Let $\bm{\delta}=(\bm{\Delta}_{1},\dots,\bm{\Delta}_{n})$ denote backdoor patterns applied to $n$ training instances in $\mathcal{D}$, and let $\mathcal{B}_{\bm{x}}$ denote the backdoor trigger added to the testing instance $\bm{x}$. Define $y_{A}=\arg\max_{y}\mathcal{P}_{\bm{\epsilon}(Z,D)}(f(\bm{x}+\mathcal{B}_{\bm{x}}+Z,\mathcal{D}+\bm{\delta}+D)=y)$, assume that when classifying a point $\mathcal{N}(\bm{x},\sigma^{2}\mathrm{I})$, the base classifier $f(\bm{x},\mathcal{D})$ assigns the most probable class $y_{A}$ with probability $P_{A}(\bm{x},\mathcal{D},\sigma)=\mathcal{P}_{\bm{\epsilon}(Z,D)}(f(\bm{x}+\mathcal{B}_{\bm{x}}+Z,\mathcal{D}+\bm{\delta}+D)=y_{A})$, and the “runner–up” class $y$ with probability $P_{B}(\bm{x},\mathcal{D},\sigma)=\max_{y\neq y_{A}}\mathcal{P}_{\bm{\epsilon}(Z,D)}(f(\bm{x}+\mathcal{B}_{\bm{x}}+Z,\mathcal{D}+\bm{\delta}+D))=y)$. Then, it is always true that $g(\bm{x}+\mathcal{B}_{\bm{x}},\mathcal{D},\sigma)=g(\bm{x}+\mathcal{B}_{\bm{x}},\mathcal{D}+\bm{\delta},\sigma)=y_{A}$ as long as the backdoor patterns $\sqrt{\sum_{i=1}^{n}\left\|\bm{\Delta}_{i}\right\|_{2}^{2}}\leq r$, where

### v2

# Cert-SSBD: Certified Backdoor Defense with Sample-Specific Smoothing Noises

@@ Abstract
Deep neural networks (DNNs) are vulnerable to backdoor attacks, where an attacker manipulates a small portion of the training data to implant hidden backdoors into the model. The compromised model behaves normally on clean samples but misclassifies backdoored samples into the attacker-specified target class, posing a significant threat to real-world DNN applications. Currently, several empirical defense methods have been proposed to mitigate backdoor attacks, but they are often bypassed by more advanced backdoor techniques. In contrast, certified defenses based on randomized smoothing have shown promise by adding random noise to training and testing samples to counteract backdoor attacks. In this paper, we reveal that existing randomized smoothing defenses implicitly assume that all samples are equidistant from the decision boundary. However, it may not hold in practice, leading to suboptimal certification performance. To address this issue, we propose a certified backdoor defense method with sample-specific smoothing noises, termed Cert-SSBD. Cert-SSBD first employs stochastic gradient ascent to optimize the noise magnitude for each sample, ensuring a sample-specific noise level that is then applied to multiple poisoned training sets to retrain several smoothed models. After that, Cert-SSBD aggregates the predictions of multiple smoothed models to generate the final robust prediction. In particular, in this case, existing certification methods become inapplicable since the optimized noise varies across different samples. To conquer this challenge, we introduce a storage-update-based certification method, which dynamically adjusts each sample’s certification region to improve certification performance. We conduct extensive experiments on multiple benchmark datasets, demonstrating the effectiveness of our proposed method. Our code is available at https://github.com/NcepuQiaoTing/Cert-SSBD.
@@ I Introduction
Figure 1: An overview of existing randomized smoothing-based certified backdoor defenses and our Cert-SSBD, providing an abstract illustration of the input space and decision boundaries. Existing methods apply fixed noise to smooth classifiers for all inputs, ignoring sample diversity, which often leads to suboptimal certification performance. In contrast, Cert-SSBD optimizes the noise, enabling the smoothing strategy to adapt to different inputs (as shown in the right figure), thereby achieving more robust certified backdoor defenses.
@@
To mitigate backdoor threats, researchers have proposed a wide range of backdoor defense strategies, including backdoor detection [64, 62, 63] and mitigation-based approaches [49, 68, 6]. However, advanced backdoor attacks [42, 47, 14] can still easily bypass existing defenses, leading to an ongoing arms race between defenders and attackers. To address this issue, some studies have proposed certified backdoor defense methods, primarily categorized into deterministic certification [32, 58, 50, 71] and probabilistic certification [55, 59]. These methods aim to provide theoretical guarantees, ensuring that the classification results of testing samples remain consistent regardless of whether the model is trained on clean or backdoor data, as long as the perturbation induced by the trigger remains within an $\ell_{p}$ norm ball of radius $r$. However, deterministic methods face scalability challenges when applied to large-scale neural networks. Consequently, probabilistic certification approaches based on randomized smoothing have emerged as a more practical alternative and have demonstrated robustness on large-scale datasets such as ImageNet [10]. Randomized smoothing was initially developed to certify robustness against adversarial examples. Its principle is to introduce random noise into the input data, ensuring that the classification results remain consistent within a specified region ($e.g.$, an $\ell_{p}$ norm neighborhood), thereby achieving robustness. Notably, pioneering studies [55, 59] showed that certified backdoor defenses based on random smoothing, which are robust against bounded backdoor patterns ($i.e.$, constrained pixel-level perturbation), can be achieved by introducing isotropic Gaussian noise into a tuple consisting of a testing instance and the training set to mitigate the impact of attacker-injected triggers, effectively neutralizing backdoor attacks during the training phase.
@@
Fortunately, the answer to the above question is affirmative. Arguably, the most direct approach is to optimize the noise at each sample by maximizing the confidence margin between the top-1 and top-2 predicted classes of the classifier ($i.e.$, the certification radius). However, the certification radius does not admit a closed-form analytical expression, which renders direct analytical optimization or deterministic gradient-based methods inapplicable. To overcome this limitation, we optimize a Monte Carlo–estimable surrogate objective that is tightly coupled with the certification radius. Inspired by the approach of [70], we adopt stochastic gradient ascent to optimize this surrogate, enabling the learning of an optimal noise level on a per-sample basis. Nevertheless, dynamically adjusting the noise during optimization inevitably alters the underlying data distribution, which increases the variance of gradient estimates and undermines optimization stability. To address this issue, we propose an advanced certified backdoor defense method with sample-specific smoothing noises, termed Cert-SSBD. In general, Cert-SSBD consists of two main stages: training and inference. In the first stage, we train multiple smoothed models using the optimized noise, which is obtained through stochastic gradient ascent to maximize the certification radius. Generally, the certification radius is computed based on the predictions of classifiers trained with fixed noise. Besides, we adopt a reparameterization technique to reduce gradient variance and enhance optimization stability. In the inference stage, we aggregate multiple smoothed classifiers trained in the first stage to generate the final smoothed prediction. However, since the optimized noise results in different noise magnitudes for each sample, existing certification methods, which typically assume a fixed noise level, are no longer directly applicable. To resolve this issue, we propose a storage-update-based certification method, which dynamically adjusts the certification region ($i.e.$, the space covered by the certification radius) for each sample. This ensures that certification regions do not overlap between different samples and that predictions remain consistent within each certified region.
@@
  • We propose a sample-specific certified backdoor defense method ($i.e.$, Cert-SSBD) to dynamically adjust the smoothing noise magnitude for different samples to optimize certification performance.
@@
  • We conduct extensive experiments on benchmark datasets to validate Cert-SSBD’s effectiveness, demonstrating its superior certification performance over existing methods.
@@ II-A Backdoor Attacks
Backdoor attacks [35, 13, 16, 17] represent an emerging threat during the DNN training phase. Based on the adversarial objective, they can be categorized into two types: all-to-one attacks, which misclassify all triggered samples into a single fixed target label and are relatively straightforward; and all-to-all attacks, which map samples to specific target classes based on their original categories, making them more complex. Besides, backdoor attacks can also be categorized based on the threat scenario into three major types: (1) poison-only attacks [19, 42], (2) training-controlled attacks [36, 39], and (3) model-modified attacks [53, 2]. Specifically, poison-only attacks restrict the adversary to modifying the training dataset; training-controlled attacks allow the adversary to fully control the training process, including both the training data and algorithms. In contrast to these approaches, model-modified attacks mainly focus on the deployment phase rather than the training phase, embedding hidden backdoors by directly modifying model weights or introducing malicious DNN modules. In this paper, we mainly focus on poison-only backdoor attacks, which represent the most classical setting and pose the broadest threat scenarios. Recently, there are also a few works exploring how to exploit backdoor attacks for positive purposes [41, 21, 60, 34, 40, 52], which is out of the scope of this paper.
@@ II-B2 Certified defenses
Existing certified defense methods can be categorized into deterministic defenses [32, 58, 50, 71], which provide guaranteed outcomes but face scalability issues, and probabilistic defenses [55, 59], ensure a ‘certified’ result with a certain probability ($e.g.$, 99.9%), where the randomness is independent of the input sample. In this work, we focus on probabilistic certified defenses.
Probabilistic certification offers better scalability. Previous methods primarily relied on intrinsic mechanisms [26, 27] or randomized smoothing techniques [55, 59] to achieve robust predictions. For example, Jia et al. [26, 27] leveraged ensemble techniques in bagging or majority voting in $k$-nearest, but these remain unsuitable for backdoor defense as they do not consider trigger size. Subsequently, Wang et al. [55] first applied randomized smoothing to backdoor defense, yet the approach lacked comprehensive evaluation and high robustness bounds. Recently, the RAB framework [59] established a theoretical foundation for provable defenses in this field (see Section II-C for more details). However, current methods implicitly assume that all samples are equidistant from the decision boundary, which may not hold in practice. This leads to suboptimal certification and highlights the urgent need for adaptive approaches that account for sample-specific characteristics.
@@ II-C Randomized Smoothing and RAB
$g(\bm{x},\sigma)\triangleq\arg\max_{y\in\mathcal{Y}}\mathcal{P}_{\bm{\epsilon}}(f(\bm{x}+\bm{\epsilon})=y),$ (1)
where $\bm{\epsilon}\sim\mathcal{N}(0,\sigma^{2}I)$. The noise level $\sigma$ is a hyperparameter that controls the trade-off between robustness and accuracy; it does not change with the input $\bm{x}$. Using the Neyman-Pearson lemma [46], Cohen et al. [8] proved that $g(\bm{x},\sigma)$ is certifiably robust to adversarial perturbations under the $\ell_{2}$ norm constraint. Define $y_{A}=\arg\max_{y}\mathcal{P}_{\bm{\epsilon}}(f(\bm{x}+\bm{\epsilon})=y)$, and assume that when classifying a perturbed input $\bm{x}+\bm{\epsilon}$, the base classifier $f$ assigns the most probable class $y_{A}$ with probability $P_{A}=\mathcal{P}_{\bm{\epsilon}}(f(\bm{x}+\bm{\epsilon})=y_{A})$, and the second most probable class $y_{B}$ with probability $P_{B}=\max_{y_{B}\neq y_{A}}\mathcal{P}_{\bm{\epsilon}}(f(\bm{x}+\bm{\epsilon})=y)$. Then, it is always true that $g(\bm{x}+\bm{\Delta},\sigma)=y_{A}$ as long as $\|\bm{\Delta}\|_{2}<r$, where the certified robust radius $r$ is given by:
$r(\bm{x},\sigma)\triangleq\frac{\sigma}{2}\left(\Phi^{-1}(P_{A}(\bm{x},\sigma))-\Phi^{-1}(P_{B}(\bm{x},\sigma))\right),$ (2)
@@
Overview of RAB [59]. Given a dataset $\mathcal{D}$ and a testing instance $\bm{x}$, the base classifier $f$ induces a predictive distribution over class labels under random perturbations: $p_{f}(y|\bm{x},\mathcal{D})\triangleq\mathcal{P}_{\bm{\epsilon}}\big(f(\bm{x},\mathcal{D})=y\big)$. The predicted label is then given by $f(\bm{x},\mathcal{D})\triangleq\arg\max_{y}p_{f}(y|\bm{x},\mathcal{D})$. A smoothed classifier $g(\bm{x},\mathcal{D},\sigma)$ returns whichever class the base classifier $f(\bm{x},\mathcal{D})$ is most likely to predict when $\bm{x}$ is perturbed by smoothing distributions $X=(Z,D)$:
@@
where $Z\sim\mathcal{N}(0,\sigma^{2}I)$ is assumed to be independent, and $D\sim\mathcal{N}(0,\sigma^{2}I)$ consists of $n$ independent and identically distributed random variables $D^{(i)}$, each added to a training instance in $\mathcal{D}$. Let $\bm{\delta}=(\bm{\Delta}_{1},\ldots,\bm{\Delta}_{n})$ denote the collection of training-set perturbations, where $\bm{\Delta}_{i}=\bm{0}$ for benign training samples and $\bm{\Delta}_{i}\neq\bm{0}$ only for poisoned training samples, and let $\mathcal{B}_{\bm{x}}$ denote the backdoor trigger added to the testing instance $\bm{x}$. Define $y_{A}=\arg\max_{y}\mathcal{P}_{\bm{\epsilon}(Z,D)}(f(\bm{x}+\mathcal{B}_{\bm{x}}+Z,\mathcal{D}+\bm{\delta}+D)=y)$, assume that when classifying a point $\mathcal{N}(\bm{x},\sigma^{2}\mathrm{I})$, the base classifier $f(\bm{x},\mathcal{D})$ assigns the most probable class $y_{A}$ with probability $P_{A}(\bm{x},\mathcal{D},\sigma)=\mathcal{P}_{\bm{\epsilon}(Z,D)}(f(\bm{x}+\mathcal{B}_{\bm{x}}+Z,\mathcal{D}+\bm{\delta}+D)=y_{A})$, and the “runner–up” class $y_{B}$ with probability $P_{B}(\bm{x},\mathcal{D},\sigma)=\max_{y_{B}\neq y_{A}}\;\mathcal{P}_{\bm{\epsilon}(Z,D)}\!\left(f(\bm{x}+\mathcal{B}_{\bm{x}}+Z,\mathcal{D}+\bm{\delta}+D)=y\right)$. Then, it is always true that $g(\bm{x}+\mathcal{B}_{\bm{x}},\mathcal{D},\sigma)=g(\bm{x}+\mathcal{B}_{\bm{x}},\mathcal{D}+\bm{\delta},\sigma)=y_{A}$ as long as the training-set perturbations satisfy $\sqrt{\sum_{i=1}^{n}\left\|\bm{\Delta}_{i}\right\|_{2}^{2}}\leq r$, where